OpenStudy (chris215):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (chris215):

cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x cos 4x + cos 2x = 1 + 1 − 2sin^2(2x) − 2sin^2(x) cos 4x + cos 2x = (1 - 2sin^2(2x)) + (1 - 2 sin^2(x) cos 4x + cos 2x = cos 4x + cos 2x is my work correct

OpenStudy (irishboy123):

yes